the length of a rectangle is three more than it's width and the perimiter of the rectangle is 58 centimeters and what are the rectangles dementions

Answer by stanbon(75887)

the length of a rectangle is three more than it's width and the perimeter of the rectangle is 58 centimeters and what are the rectangles demensions
------
L = W+3
Perimeter = 2(L + W)
58 = 2(W+3 + W)
29 = 2W+3
2W = 26
W = 13 (width)
L = W+3 = 16 (length)
=========================
Cheers,
Stan H.
===============